You are bearish on Telecom and decide to sell short 280 shares at the current market price of $100 per share. a. How much in cash or securities must you put into your brokerage account if the brokers initial margin requirement is 50% of the value of the short position? (Round your answer to the nearest whole number) or to be put into b. How high can the price of the stock go before you get a margin call if the maintenance margin is 30% of the value of the short position? (Round your answer to 2 decimal places.) or
Answer #1

Please provide rating..

Answer a)
Number of share shorted = 280
Price of share = 100
Total value = 28000
Cash or securities required in brokerage account = 28000*50% 14000
Answer b)
Maintenance margin = 30%
Lets assume that higher price at which margin call will be receive d= X
=(28000+14000-100*X)/100*X = 0.3
=42000-100X = 30X
130X = 42000
x     323.08
Therefore higher price at which will have margin call =     323.08 or higher